OpenStudy (anonymous):

Solve the system of equations given as x^2 +y^2 =16 y-4 = -x^2

OpenStudy (anonymous):

We have x^2=4-y , from the second equation.So we replace it to the first one : 4-y+y^2=16 @Romwil

OpenStudy (campbell_st):

well you have a circle with a centre at the origin and readius 4 and a concave down parabola with a vertex at (0, 4)

OpenStudy (campbell_st):

so there are 3 points of intersection...

OpenStudy (anonymous):

Refer to the attachment generated with the Mathematica 9 program.

OpenStudy (sepeario):

props to him ^^ he has taken the time to graphically solve it with a calculator, rather than completely algebraically :)